# Capture of field stars by dark substructures

### v1

@@ 3 Statistical Experiments
§3.2 injects a static substructure potential $\Phi_{\bullet}$ in a pre-existing field of particles in dynamical equilibrium within $\Phi^{\rm dSph}$. This set-up is similar to the analytical conditions adopted in §2.1, and we can therefore anticipate a close match between theoretical predictions and the numerical results. Interestingly, we will see that the dynamics of field particles trapped in $\Phi_{\bullet}$ can be broadly separated in two families: (i) permanent captures, which become bound immediately after the substructure is placed in the field at $t=0$, and remain bound for arbitrarily long times, and (ii) temporary captures, which as the name indicates only remain bound for a finite amount of time.
@@ 3.1 Initial conditions & set-up
For illustration, we choose a substructure mass $M_{\bullet}=10^{6}\,{\rm M}_{\odot}$. To estimate the thermal critical radius we adopt $\sigma=\langle v^{2}\rangle^{1/2}/3=12.2\,{\rm km\,s^{-1}}$ in Equation (6), which returns $r_{\epsilon}=18.9\,{\rm pc}$. Note that this size is comparable to that of the extended stellar clusters discussed in the Introduction. We consider a ‘compact’ substructure with $\kappa=+0.8$, which translates into a scale radius $c_{\bullet}=(1-\kappa)\,r_{\epsilon}=3.78\,{\rm pc}$ (notice that this value is smaller than the typical scale radius of field haloes with a vitial mass $M_{\rm vir}(z=0)\sim 10^{6}M_{\odot}$ in the Aquarius simulation, see Fig. 2). The time-scale needed to cross the over-density size by a random field particle is much shorter than the orbital time, $T_{\epsilon}=r_{\epsilon}/\sigma=18.9\,{\rm pc}/(12.2\,{\rm km\,s^{-1}})\simeq 1.5\,{\rm Myr}$. As a result, in these models the population of captured particles reaches steady state quickly.
@@
Figure 3: Distribution of capture times ($t_{c}=t-t_{\rm bound}$) defined as the time span since a particle becomes bound ($t_{\rm bound}$), i.e. the energy $E$ flips from positive to negative, to the current simulation time ($t$). Time is measured in units Smoluchowski’s time scale ($T$). Here, we only include particles within a volume size equal to the local tidal radius, $r_{\rm out}=r_{t}$ (see text). Notice the presence of a substantial number of particles that become bound at $t\approx 0$, which result in a a strong peak in the distribution at $t_{c}\approx t$ shifting towards the right as the integration time increases.
@@ 3.2 Static substructures potential on circular orbits
In this Section we place a substructure potential $\Phi_{\bullet}$ at a galactocentric radius $R_{\bullet}$ with a circular velocity $V_{\bullet}=V_{c}(R_{\bullet})$ in a sea of tracer particles with an extended profile (see §3.1) in dynamical equilibrium. By chance, at the initial snapshot ($t=0$) a number of tracer particles are found within a volume element $V=4\pi r^{3}_{\rm out}/3$ centred at $\Phi_{\bullet}$ with a relative velocity below the escape velocity, $v<v_{e}$, which implies a negative energy, $E<0$. In what follows, we refer to these particles as ‘immediate captures’ in order to distinguish them from ‘chaotic 3-body captures’ (e.g. Petit & Henon 1986) taking place during the dynamical integration of their orbits as a result of the interplay between the tracer particle and the substructure and host galaxy potentials.
@@ 3.3 Growing substructures on a circular orbit
Fig. 7 shows the distribution of capture times, $t_{c}=t-t_{\rm bound}$, where $t_{\rm bound}$ is the integration time at which a particle becomes bound, in models with three different mass-growth rates, $\tau/T=0,1$ and $3$. Orange lines correspond to the same distribution plotted in Fig. 3 for static models ($\tau=0)$. Comparison between the panels from top to bottom shows the emergence of prominent spike of particles with $t_{c}\approx t$, which is the telltale of the build-up of a permanent captures trapped in the time-evolving substructure potential, $\Phi_{\bullet}(t)$. Focusing on the models with a slow growth rate ($\tau=3\,T$, red lines) we find that the spike is absent early on during the simulation ($t/T=5$, top panel), and that it only starts to arise at $t/T\gtrsim 20$ (second panel from the top). At the end of the simulation at $t=50\,T$ (third panel from the top), we find that all three models show a clear excess of captures with long capture times, $t_{c}\sim t$, with the peak of the distribution broadening up for models with slow growth rates ($\tau>0$). This suggests that while in a static potential ($\tau=0$) permanent captures are trapped immediately at $t=0$, this is not the case in substructures with a mass growth (22), which capture particles onto permanent orbits over an extended period of time, $t\lesssim 2\tau$.
@@ 3.4 Accreted substructures on an eccentric orbit
Figure 11: Number of bound particles located within a volume size $r_{\rm out}=0.3\,{\rm kpc}$ as a function of time for the models plotted in Fig. 10. Black-dashed lines show the theoretical prediction from Equation (2), with the time-dependent number density and velocity dispersion of field particles measured at the location of the substructure, $n(t)=n[R(t)]$ and $\sigma(t)=\sigma[R(t)]$. Pericentric passages are marked with vertical dotted lines.
@@
The rate at which field particles are captured in these two models is also strikingly different. Fig. 11 shows the number of bound particles within a volume size $r_{\rm out}=300\,{\rm pc}$ (marked with black circles in Fig. 10 for reference) as a function time. The top panel shows that the substructure potential orbiting in the extended galaxy contains $N_{b}\sim 3\times 10^{3}$ at $t=0$. These are immediate captures by definition. As the substructure falls into the inner regions of the galaxy, the number of bound particles grows slightly, suffering a significant increase after the first pericentric passage. At the last snapshot of the simulation, the number of bound particles has increased by a factor $\sim 10$ up to $N_{b}(t_{f})\sim 4\times 10^{4}$. Remarkably, the theoretical expectation given by Equation (2) (black-dashed line) barely changes along the orbital phase in spite of the large variation of galactocentric distances and velocities. This result reflects the counter balance between the variation of the mean phase-space density of field particles along the orbit, $Q(R)=n(R)/\sigma^{3}(R)$, and the exponential suppression of captures at high velocities, $\xi=\exp[-V^{2}_{\bullet}/(2\sigma^{2}(R))]$. As the substructure plunges into the inner regions of the galaxy the orbital velocity increases, making capture less efficient. Simultaneously, it encounters field particles with higher phase-space densities, with both effects nearly balancing out in Equation (2). The reverse process takes place from pericentre to apocentre.
@@
In contrast, the substructure orbiting in the truncated galaxy does not capture field particles until it is close to its first orbital pericentre at $\Omega\,t\simeq 1.2$. Bottom panel of Fig. 11 shows a rapid increase in the population of bound particles, which grows $\sim 3$ orders of magnitude after the pericentric passage. A large fraction of these particles are quickly lost to galactic tides, such that after its first apogalacticon the number of bound particles plateaus at $N_{b}\sim 10^{2}$. As it reaches its second pericentre, it grows up again to $N_{b}\sim 5\times 10^{3}$, falling back to $N_{b}\sim 10^{2}$ at its second apocentre. Comparison with the theoretical expectation (black-dashed lines) shows that Equation (2) cannot accurately describe this behaviour. There are two main reasons for the mismatch. In the first place, the statistical theory derived in §2.1 relies on the local approximation, which assumes that the properties of the galactic field do not vary strongly as a function of distance from the substructure. This is clearly not the case in this particular galaxy model, which is strongly truncated beyond $R\gtrsim 600\,{\rm pc}$. Second, it also assumes that the properties of the field are time-invariant, and that the population of captured particles is in steady state at all times. None of these conditions apply to the these particular models, which are strongly time dependent. The latter approximation is particularly poor after the first pericentric passage. Here, the numerical experiments show that the substructure has captured $N_{b}\sim 10^{3}$ particles after the first orbital revolution. In contrast, Equation (2) predicts $N_{b}=0$ because the density of field particles at this location is $n(R_{\rm apo})=0$ (see bottom-left panel of Fig. 10). This mismatch repeats again at the second apocentre. Clearly, we observe field particles in the vicinity of the substructure in the bottom-right panel of Fig. 10, which are not predicted by Equation (2. These particles are either still bound to the substructure (red dots), or were ‘scooped’ from the inner regions of the galaxy and released from the substructure by galactic tides (black dots).
@@
The substructure orbiting in an extended galaxy contains a population of captured field particles with a profile that roughly follows Equation (4). This is a remarkable result considering the strong strong temporal evolution of the field along the substructure orbit, and that the statistical theory is derived for static systems. The good agreement suggests the (local) steady-state expressions are accurate at leading order insofar as the orbital period around the host galaxy is longer than the internal crossing time of the substructure. In our models, the crossing time ($T_{\epsilon}=r_{\epsilon}/\langle v^{2}\rangle^{1/2}\simeq 1.5\,{\rm Myr}$) is much shorter than their orbital time in the dSph potential ($t_{\rm dyn}=R_{\bullet}/V_{\bullet}\sim 30\,{\rm Myr}$), which suggests that the steady-state equations are reasonably accurate far from orbital pericentre.
In contrast, field particles bound to the substructure orbiting in an extended galaxy show a profile that falls off more steeply than predicted by Equation (4) at large distances, whereas at small distances, $r\ll c_{\bullet}$, it matches the profile of the extended galaxy model.
In general, the numerical experiments shown in this Section indicate that the distribution of stars captured by moving substructures depend on whether these objects are form in a pre-existing sea of field particles, or are accreted from outside the galaxy.
@@ 4.2 Theory: simplifications and current uncertainties
The statistical experiments in §3 adopt substructure models with masses that are either constant, or grow with time. The latter models can be applied for example to study the population of field particles trapped around intermediate mass black holes (IMBHs) that grow in disks surrounding supermassive black holes (e.g. McKernan et al. 2012). However, our models do not cover the case of self-gravitating subhaloes orbiting in a parent halo, which experience tidal stripping and mass loss after each pericentre passage (e.g. Peñarrubia et al. 2010). This shortcoming may be addressed by running live $N$-body simulations where substructures are modelled as self-gravitating objects. Unfortunately, the numerical tools required to explore this scenario are more complex and computationally expensive than the ones used for this work. The results enclosed in this paper suggest that the analytical expressions outlined in §2 can be applied to time-dependent substructures, $\Phi_{\bullet}(t)$, insofar as its time-evolution is not impulsive, i.e.t it occurs on time scales $|\dot{\Phi}_{\bullet}/\Phi_{\bullet}|^{-1}\gtrsim T$, and the local approximation is reasonably accurate. Under these conditions, the population of bound field stars can be assumed to be in steady state in a time-varying potential $\Phi_{\bullet}(t)$, see Fig. 8. For substructures moving on eccentric orbits around extended galaxies, numerical experiments in §3.4) show that the steady-state approximation may be reasonably accurate away from orbital pericentre. In galaxies with steep density profiles, it is the local approximation that fails away from orbital pericentre.
@@ 4.3 Dark Matter particle constraints
Detecting field stars trapped in a dark subhalo with a mass of $M_{\bullet}\simeq 10^{6}\,{\rm M}_{\odot}\gtrsim m_{\rm hm}$ would imply a half-mass mode of $m_{\rm hm}\gtrsim 18\,{\rm keV}$, significantly tightening existing constraints from gravitational lensing ($m_{\rm DM}\gtrsim 5\,{\rm keV}$; Gilman et al. 2020), Lyman-alpha forest ($m_{\rm DM}\gtrsim 3\,{\rm keV}$; Villasenor et al. 2022), or the combination of strong gravitational lensing, the Ly-$\alpha$ forest, and the number of luminous satellites in the Milky Way, which put a lower particle mass limit of $m_{\rm DM}\gtrsim 6\,{\rm keV}$ (Enzi et al. 2021). Detecting objects with masses $M_{\bullet}\simeq 10^{5}\,{\rm M}_{\odot}\gtrsim m_{\rm hm}$ would push the constraints up to $m_{\rm hm}\gtrsim 37\,{\rm keV}$.
@@
In addition, the internal structure of these systems may also reveal whether DM particles experience self-interactions. For example, dark substructures with $M_{\bullet}\sim 10^{6}\,{\rm M}_{\odot}$ must be sufficiently compact, $c_{\bullet}\lesssim 20\,{\rm pc}$, in order to generate a stellar overdensity of field stars in a Fornax-like dSph (see middle panel of Fig. 2). This condition translates into a characteristic density $\rho_{\bullet}=M_{\bullet}/(2\pi c_{\bullet}^{3})\gtrsim 20\,{\rm M}_{\odot}\,{\rm pc}^{-3}$, which is higher than CDM subhaloes with similar masses. More precisely, field halos with a mass of $M_{200}=10^{6}\,{\rm M}_{\odot}$ at redshift $z=0$ have a mean peak velocity radius of $r_{\rm max}\sim 100\,{\rm pc}$. Equating this radius to the substructure scale radius, $c_{\bullet}=r_{\rm max}$, yields a considerably lower characteristic density of $\rho_{\bullet}\sim 0.1\,{\rm M}_{\odot}\,{\rm pc}^{-3}$ (Ludlow et al. 2016). On the one hand, the existence of DM sub-subhaloes with abnormally-high densities may point to interactions in the dark sector (e.g. Kahlhoefer et al. 2019). On the other, it is not immediately clear whether such a high density would necessarily be in conflict with CDM predictions. First, because sub-subhalos in satellite galaxies are expected to be tidally processed, and more concentrated than field halos of the same mass (e.g. Peñarrubia et al. 2010; Errani & Navarro 2021). And second, because it is possible that the detection of dense sub-subhaloes may be due to observational biases. Indeed, compact substructures produce higher density enhancements in the field (see §2), which can therefore be more easily detected as localized overdensities ($\delta_{\star}>1$). In this picture, anomalous objects like F6 may be sampling the high-density, low-probability tail of the sub-subhalo population. That is, they would represent the proverbial ‘tip of the iceberg’, pointing to the presence of a much larger population of diffuse ($\delta_{\star}<1$) substructures with a power-law luminosity function (15). There are several ways to test this prediction, e.g. by modelling statistical fluctuations of number counts in photometric surveys (e.g. Scheuer 1957), or by searching for clumps with low-velocity dispersion and/or significant velocity offsets in kinematic surveys of dSphs (e.g. Pace et al. 2014).

### v2

@@ 3 Statistical Experiments
§3.2 injects a static substructure potential $\Phi_{\bullet}$ in a pre-existing field of particles in dynamical equilibrium within $\Phi^{\rm dSph}$. This set-up is similar to the analytical conditions adopted in §2.1, and we can therefore anticipate a close match between theoretical predictions and the numerical results. Interestingly, we will see that the dynamics of field particles trapped in $\Phi_{\bullet}$ can be broadly separated in two families: (i) permanent captures, which become bound immediately after the substructure is placed in the field at $t=0$, and remain bound for indefinitely long times, and (ii) temporary captures, which as the name indicates only remain bound for a finite amount of time.
@@ 3.1 Initial conditions & set-up
For illustration, we choose a substructure mass $M_{\bullet}=10^{6}\,{\rm M}_{\odot}$. To estimate the thermal critical radius we adopt $\sigma=\langle v^{2}\rangle^{1/2}/3=12.2\,{\rm km\,s^{-1}}$ in Equation (9), which returns $r_{\epsilon}=18.9\,{\rm pc}$. Note that this size is comparable to that of the extended stellar clusters discussed in the Introduction. We consider a ‘compact’ substructure with $\kappa=+0.8$, which translates into a scale radius $c_{\bullet}=(1-\kappa)\,r_{\epsilon}=3.78\,{\rm pc}$ (notice that this value is smaller than the typical scale radius of field haloes with a vitial mass $M_{\rm vir}(z=0)\sim 10^{6}M_{\odot}$ in the Aquarius simulation, see Fig. 2). The time-scale needed to cross the over-density size by a random field particle is much shorter than the orbital time, $T(r_{\epsilon})=r_{\epsilon}/\sigma=18.9\,{\rm pc}/(12.2\,{\rm km\,s^{-1}})\simeq 1.5\,{\rm Myr}$. As a result, in these models the population of captured particles reaches steady state quickly.
@@
Figure 3: Distribution of capture times ($t_{c}=t-t_{\rm bound}$) defined as the time spanned since a particle becomes bound ($t_{\rm bound}$), i.e. the energy $E$ flips from positive to negative, until the present integration time ($t$). Time is measured in units Smoluchowski’s time scale ($T$) for convenience. Here, we only include particles within a volume size equal to the local tidal radius, $r_{\rm out}=r_{t}$ (see text). Notice the presence of a substantial number of particles that become bound at $t\approx 0$, which result in a a strong peak in the distribution at $t_{c}\approx t$ shifting towards the right as the integration time increases.
@@ 3.2 Static substructures potential on circular orbits
In this Section we place a substructure potential $\Phi_{\bullet}$ at a galactocentric radius $R_{\bullet}$ with a circular velocity $V_{\bullet}=V_{c}(R_{\bullet})$ in a sea of tracer particles with an extended profile (see §3.1) in dynamical equilibrium. By chance, at the initial snapshot ($t=0$) a number of tracer particles are found within a volume element $V=4\pi r^{3}_{\rm out}/3$ centred at $\Phi_{\bullet}$ with a relative velocity below the escape velocity, $v<v_{e}$, which implies a negative energy, $E<0$. In what follows, we refer to these particles as ‘immediate captures’ in order to distinguish them from ‘chaotic 3-body captures’ (e.g. Petit & Henon 1986), which take place during the dynamical integration of their orbits as a result of the interplay between the tracer particle and the substructure and host galaxy potentials.
@@ 3.3 Growing substructures on a circular orbit
Fig. 7 shows the distribution of capture times, $t_{c}=t-t_{\rm bound}$, where $t_{\rm bound}$ is the integration time at which a particle becomes bound, in models with three different mass-growth rates, $\tau/T=0,1$ and $3$. Orange lines correspond to the same distribution plotted in Fig. 3 for static models ($\tau=0)$. Comparison between the panels from top to bottom shows the emergence of prominent spike of particles with $t_{c}\approx t$, which is the telltale of the build-up of a permanent captures trapped in the time-evolving substructure potential, $\Phi_{\bullet}(t)$. Focusing on the models with a slow growth rate ($\tau=3\,T$, red lines) we find that the spike is absent early on during the simulation ($t/T=5$, top panel), and that it only starts to arise at $t/T\gtrsim 20$ (second panel from the top). At the end of the simulation at $t=50\,T$ (third panel from the top), we find that all three models show a clear excess of captures with long capture times, $t_{c}\sim t$, with the peak of the distribution broadening up for models with slow growth rates ($\tau>0$). This suggests that while in a static potential ($\tau=0$) permanent captures are trapped immediately at the start of the experiment, $t_{\rm bound}=0$, this is not the case in substructures with a mass growth (21), which capture particles onto permanent orbits over an extended period of time, $t_{\rm bound}\lesssim 2\tau$.
@@ 3.4 Accreted substructures on an eccentric orbit
Figure 11: Number of bound particles located within a volume size $r_{\rm out}=0.3\,{\rm kpc}$ as a function of time for the models plotted in Fig. 10. Dotted lines show the total (bound plus unbound) number of field particles within the volume, $r<r_{\rm out}$. Black-dashed lines show the theoretical prediction from Equation (2), with a time-dependent number density and velocity dispersion of field particles measured at the location of the substructure, $n(t)=n[R(t)]$ and $\sigma(t)=\sigma[R(t)]$. Pericentric passages are marked with vertical dotted lines.
@@
The rate at which field particles are captured in these two models is also strikingly different. Solid lines in Fig. 11 shows the number of bound particles within a volume size $r_{\rm out}=300\,{\rm pc}$ (marked with black circles in Fig. 10 for reference) as a function time. In the top panel we can see that the substructure orbiting in the extended galaxy contains $N_{b}\sim 3\times 10^{3}$ at $t=0$. By definition, these are immediate captures. As the substructure falls into the inner regions of the galaxy, the total number of field particles within the volume (dotted lines) increases systematically until pericentre is reached at $\Omega\,t\simeq 1.4$. After the pericentre passage, the number of bound particles (solid lines) grows slightly, reaching a plateau at apocentre, $\Omega\,t\simeq 2.8$. This situation repeats throughtout the next orbital revolution. At the last snapshot of the simulation, the number of bound particles is $N_{b}(t_{f})\sim 4\times 10^{4}$, which is a factor $\sim 10$ increases with respect to the value at $t=0$. Remarkably, the theoretical expectation given by Equation (2) (black-dashed line) barely changes as a function of orbital phase in spite of the large variation of galactocentric distances and velocities. This reflects the counter balance between the variation of the mean phase-space density of field particles along the orbit, $Q(R)=n(R)/\sigma^{3}(R)$, and the exponential suppression of captures at high velocities, $\xi=\exp[-V^{2}_{\bullet}/(2\sigma^{2}(R))]$. As the substructure plunges into the inner regions of the galaxy the orbital velocity increases, making capture less efficient. Simultaneously, the number of field particles around the substructure grows, with both effects nearly balancing out in Equation (2). The reverse process takes place from pericentre to apocentre.
@@
The substructure model orbiting within a truncated galaxy experiences a very different evolution. In this case, bottom panel of Fig. 11 shows that capture does not occur until the substructure is close to its first orbital pericentre. As expected, the number of immediate captures is zero by construction. After each pericentric passage, we observe a very rapid increase in the population of bound particles, whose size can grow by $\sim 3$ orders of magnitude. However, a large fraction of these particles are quickly lost to galactic tides. At first apogalacticon, the number of bound particles has stabilized at $N_{b}\sim 10^{2}$. As it reaches its second pericentre, it grows again up to $N_{b}\sim 5\times 10^{3}$, falling back to $N_{b}(t_{f})\sim 10^{2}$ at its second apocentre. Comparison with the theoretical expectation (black-dashed lines) shows that Equation (2) cannot accurately describe this behaviour. There are two main reasons for the mismatch. In the first place, the statistical theory derived in §2.1 relies on the local approximation, which assumes that the properties of the galactic field do not vary strongly as a function of distance from the substructure. This is clearly not the case in this particular galaxy model, which is strongly truncated beyond $R\gtrsim 600\,{\rm pc}$. Furthermore, it assumes that the properties of the field are time-invariant, and that the population of captured particles is in steady state at all times. None of these conditions apply to the the model shown in the bottom panel. The latter approximation is particularly poor after the first pericentric passage. Here, Equation (2) predicts $N_{b}=0$ because the density of field particles at this location is $n(R_{\rm apo})=0$ (see bottom-left panel of Fig. 10). In contrast, our numerical experiment shows that the substructure has captured $N_{b}\sim 10^{3}$ particles after the first orbital revolution. This mismatch arises again at the second apocentre, which corresponds to the final snapshot of the simulation. Here, all particles within the volume under observation (dotted line) are bound to the substructure (solid line). It is worth stressing that region of the galaxy was empty from field particles at $t=0$. Looking the bottom-right panel of Fig. 10, we find field particles beyond the original truncation radius of the galaxy are either still bound to the substructure (red dots), or were ‘scooped’ from the inner regions and released back to the galaxy by tidal forces (black dots).
@@
Particles captured by the substructure model orbiting in an extended galaxy exhibit a profile shape that roughly scales as $(1+r/c_{\bullet})^{-3/2}$, which matches Equation (6). This is a remarkable result considering the strong strong temporal evolution of the field along the substructure orbit, and that the statistical theory is derived for static systems. However, the normalization of the profile is not so well matched by the analytical formula. In particular, Fig. 11 shows that Equation (6) tends to underpredict the number of field particles bound to the substructure at all points of the orbit.
This mismatch is considerably stronger in substructures orbiting a in a truncated galaxy. In these objects, captured field particles show a profile that falls off more steeply than predicted by Equation (6) at large distances, $r\gtrsim c_{\bullet}$, whereas at small distances, $r\ll c_{\bullet}$, it converges to the profile of the extended galaxy model. The different behaviour may be traced back to the lack of permanent captures in the truncated galaxy models, which tend to smooth out the overall density enhancement profile (see Fig. 6).
In general, the numerical experiments in this Section indicate that the distribution of field stars captured by substructures moving on eccentric orbits vary as a function of time in different ways depending on whether these objects are form in a pre-existing sea of field particles, or are accreted from outside the galaxy.
@@ 4.2 Theory: simplifications and current uncertainties
The statistical experiments in §3 adopt substructure models with masses that are either constant, or grow with time. The latter models can be applied for example to study the population of field particles trapped around intermediate mass black holes (IMBHs) that grow in disks surrounding supermassive black holes (e.g. McKernan et al. 2012). However, our models do not cover the case of self-gravitating subhaloes orbiting in a parent halo, which experience tidal stripping and mass loss after each pericentre passage (e.g. Peñarrubia et al. 2010). This shortcoming may be addressed by running live $N$-body simulations where substructures are modelled as self-gravitating objects. Unfortunately, the numerical tools required to explore this scenario are more complex and computationally expensive than the ones used for this work. The results enclosed in this paper suggest that the analytical expressions outlined in §2 can be applied to time-dependent substructures, $\Phi_{\bullet}(t)$, insofar as its time-evolution is not impulsive, i.e.t it occurs on time scales $|\dot{\Phi}_{\bullet}/\Phi_{\bullet}|^{-1}\gtrsim T$, and the local approximation is reasonably accurate. Under these conditions, the population of bound field stars can be assumed to be in steady state in a time-varying potential $\Phi_{\bullet}(t)$, see Fig. 8. For substructures moving on eccentric orbits around extended galaxy models, numerical experiments in §3.4) show that the steady-state approximation may be reasonably accurate away from orbital pericentre. In truncated galaxy models, it is the local approximation that fails away from pericentre (see Fig. 10).
@@ 4.3 Dark Matter particle constraints
Detecting field stars trapped in a dark subhalo with a mass of $M_{\bullet}\simeq 10^{6}\,{\rm M}_{\odot}\gtrsim m_{\rm hm}$ would imply a half-mass mode of $m_{\rm hm}\gtrsim 18\,{\rm keV}$, significantly tightening existing constraints from gravitational lensing ($m_{\rm DM}\gtrsim 5\,{\rm keV}$; Gilman et al. 2020), Lyman-alpha forest ($m_{\rm DM}\gtrsim 3\,{\rm keV}$; Villasenor et al. 2022), or the combination of strong gravitational lensing, the Ly-$\alpha$ forest, the number of luminous satellites in the Milky Way, which put a lower particle mass limit of $m_{\rm DM}\gtrsim 6\,{\rm keV}$ (Enzi et al. 2021; Nadler et al. 2021), and $m_{\rm DM}>9.7$ keV when the Milky Way satellite population is combined with strong-lensing flux ratio statistics (Nadler et al. 2021). Detecting objects with masses $M_{\bullet}\simeq 10^{5}\,{\rm M}_{\odot}\gtrsim m_{\rm hm}$ would push the constraints up to $m_{\rm hm}\gtrsim 37\,{\rm keV}$.
@@
In addition, the internal structure of these systems may also reveal whether DM particles experience self-interactions. For example, dark substructures with $M_{\bullet}\sim 10^{6}\,{\rm M}_{\odot}$ must be sufficiently compact, $c_{\bullet}\lesssim 20\,{\rm pc}$, in order to generate a stellar overdensity of field stars in a Fornax-like dSph (see middle panel of Fig. 2). This condition translates into a characteristic density $\rho_{\bullet}=M_{\bullet}/(2\pi c_{\bullet}^{3})\gtrsim 20\,{\rm M}_{\odot}\,{\rm pc}^{-3}$, which is higher than CDM subhaloes with similar masses. More precisely, field halos with a mass of $M_{200}=10^{6}\,{\rm M}_{\odot}$ at redshift $z=0$ have a mean peak velocity radius of $r_{\rm max}\sim 100\,{\rm pc}$. Equating this radius to the substructure scale radius, $c_{\bullet}=r_{\rm max}$, yields a considerably lower characteristic density of $\rho_{\bullet}\sim 0.1\,{\rm M}_{\odot}\,{\rm pc}^{-3}$ (Ludlow et al. 2016).
On the one hand, the existence of DM sub-subhaloes with abnormally-high densities may point to interactions in the dark sector (e.g. Kahlhoefer et al. 2019). On the other, it is not immediately clear whether such a high density would necessarily be in conflict with CDM predictions. First, because sub-subhalos in satellite galaxies are expected to be tidally processed, and more concentrated than field halos of the same mass (e.g. Peñarrubia et al. 2010; Errani & Navarro 2021). And second, because it is possible that the detection of dense sub-subhaloes may be due to observational biases. Indeed, compact substructures produce higher density enhancements in the field (see §2), which can therefore be more easily detected as localized overdensities ($\delta_{\star}>1$). In this picture, anomalous objects like F6 may be sampling the high-density, low-probability tail of the sub-subhalo population. That is, they would represent the proverbial ‘tip of the iceberg’, pointing to the presence of a much larger population of diffuse ($\delta_{\star}<1$) substructures with a power-law luminosity function (14). There are several ways to test this prediction, e.g. by modelling statistical fluctuations of number counts in photometric surveys (e.g. Scheuer 1957), or by searching for clumps with low-velocity dispersion and/or significant velocity offsets in kinematic surveys of dSphs (e.g. Pace et al. 2014).